Consider the following reaction: CO(g)+H2O(g)⇌CO2(g)+H2(g) Kp=0.0611 at 2000 K A reaction mixture initially contains a CO partial pressure of 1380 torr and a H2O partial pressure of 1770 torr at 2000 K. Calculate the equilibrium partial pressure of CO2.
Answer #1
CO(g) H2O(g) \rightleftharpoons CO2(g) H2(g)
Initial (torr) 1380 1770 0 0
Change (torr) -x -x +x +x
Equilibrium (torr) 1380-x 1770-x x x

Kp = (pCO2)(pH2) / (pCO * pH2O)

0.0611 = x2 / (1380-x) (1770-x)

0.0611 (2442600 - 3150x + x2) = x2

149242.86 - 192.465x + 0.0611x2 = x2

0.9389x2 + 192.465x - 149242.86 = 0

x = 309.16

Thus, at equilibrium partial pressure of CO2 = 309.2 torr